# Ion Mix Can Invert Centrifugal Traps

### v1

@@ Abstract
In centrifugal plasma traps, electric fields naturally arise in the direction parallel to the magnetic field in order to ensure quasineutrality. These electric fields are sensitive to the particular mix of ion species present. Changing the composition of the plasma can mix or demix different ion species, improve confinement of certain ions, and can even expel certain species from the trap altogether. This makes it possible to generate an “inverted centrifugal” end-plug with very promising properties.
Introduction. It is well-known that mirror traps tend to develop electric fields parallel to the magnetic field lines. In the absence of these fields, ions are better-confined than electrons. In non-rotating or slowly-rotating mirrors, this is mainly because of the difference in the pitch-angle scattering rates, and in rapidly-rotating centrifugal mirrors, this is mainly because centrifugal confinement acts more effectively on more massive species. In either case, electric fields arise to enforce local quasineutrality. These are sometimes called ambipolar fields.
@@
These ideas were developed for conventional (that is, non-centrifugal) mirrors. For centrifugal mirrors, the possibilities associated with these ambipolar fields have been relatively understudied. In this paper, we will consider the effects of ion species mix on ambipolar potentials in centrifugal mirrors. These mix effects lead to surprising outcomes. Adding more of one species can confine or deconfine other populations, or can lead to stratification effects in which different species are localized at different locations along field lines. Even more surprising is that mix effects can also be used to establish a highly advantageous centrifugal end-plug, analogous to the end-plugs used in conventional mirrors but relying on a different set of physical mechanisms. These effects suggest exciting new possibilities for centrifugal mirror fusion, among other applications.
@@
Here $\Delta\phi$ is the electrical potential, $\Delta\varphi_{c}$ is the centrifugal potential as it would be experienced by a proton, $q_{s}$ is the charge, $\mu_{s}$ is the particle mass (normalized by the proton mass), and we have taken the temperature $T_{s}$ to be constant along the field line. $n_{s0}$ is a normalization constant. $\Delta\phi$ and $\Delta\varphi_{c}$ are both defined relative to some reference point on the field line. In the quasineutral limit,
@@
A well in $\Delta\varphi_{c}$ always confines species $a$ and a barrier in $\Delta\varphi_{c}$ always repels species $a$; the electrostatic field simply acts to reduce the efficacy of these wells or barriers by a factor of $1+Z_{a}(T_{e}/T_{a})$.
@@
Figure 1: Different distributions of protons and boron-11 along a field line. As the relative density of the protons drops, the self-consistent ambipolar fields push them further outwards (the top panel has the highest boron concentration; the bottom panel has the highest proton concentration). This example takes the ion and electron temperatures to be equal. Note that the overall normalization is arbitrary, so we could imagine the different cases taking place at fixed total ion charge or fixed total ion number. $n_{p0}$ and $n_{B0}$ are defined as the proton and boron densities at $\varphi_{c}=0$.
@@
Consider, then, another example: a deuterium-tritium plasma that is “doped” with a smaller fraction of protons. Protons, with their high charge-to-mass ratio, tend to reduce the ambipolar fields wherever they are found. This is illustrated in Fig. 2, where the introduction of a small population of protons substantially improves the confinement of deuterium and tritium in a centrifugal well. The example in the figure introduces a relatively small proton fraction; the size of the effect on deuterium and tritium confinement depends on the size of the proton population.
So, should deuterium-tritium centrifugal traps operate with a small fraction of protons? The answer depends on which factors limit the performance of the trap. The introduction of protons allows the deuterium and tritium to be confined at lower rotation speeds, which would suggest that a pressure-limited device could operate at higher plasma densities (since rotation speeds strongly affect the pressure limits in these devices [11]). However, the protons themselves are relatively poorly confined in these scenarios. This downside might be mitigated using the centrifugal end-plugs proposed later in this paper.
@@
The same physical mechanisms discussed in the two examples above can also be applied to a wide variety of other problems. One of these is the physics of heavy impurities. Because heavy impurities are often not fully ionized, their charge-to-mass ratio can be particularly low. This means that impurity accumulation has the potential to worsen the longitudinal confinement of all other species. Another straightforward application is plasma mass filters, which are designed to separate an input of mixed ions according to species [12, 13, 14, 15, 16, 17, 18, 19]. If the composition of a centrifugal trap can be tuned to confine or deconfine different species according to mass and charge, the configuration could be used for separations.
@@
The confinement generally gets worse if other ions are placed in the well alongside the protons. However, suppose instead that the protons are placed in a region between two centrifugal wells, each populated with some other, lower charge-to-mass ratio ion – say, 7Li. In the limit where the centrifugal wells are populated entirely with lithium, the protons are repelled from the wells, seeing a total potential
@@
In other words, counter-intuitively, for a given well depth $\Delta\varphi_{c}$, it is more efficient for the purposes of proton confinement to set up a barrier full of some other, lower-$Z/m$ species than it is to use that same $\Delta\varphi_{c}$ as a well to directly trap the protons. This is illustrated in Figure 3. One could make an even more effective trap by placing the protons in a centrifugal well of their own, then flanking that well with end-cells full of another species.
@@
Another possibility is to use partially ionized species in the end-cell. For partially ionized species, the charge-to-mass ratio can be very small, so an end-cell populated by a partially ionized species can in principle produce a very strong barrier. For higher-temperature applications, there is the challenge of how to maintain a lower ionization state (and the radiative losses normally associated with partially ionized species like tungsten in high-temperature plasmas). A degree of thermal isolation between the main cell and the end-cells naturally arises if the end-cells produce potential barriers to repel main-cell ions, but the limits on how easily a temperature difference might be maintained are left as an open question. The “thermal barrier” concept for the end-cells of conventional (non-centrifugal) mirrors does rely on maintaining a temperature difference between end-cell and main-cell, and least for the electron population [1, 7]. However, it is substantially easier to maintain a difference in electron temperatures than in ion temperatures.
@@
The model used here comes with a number of simplifications. First, and perhaps most importantly: we assume that each species is Gibbs-distributed along field lines. This assumption is examined in further detail in the Appendix. Kinetically, it is equivalent to the assumption that the velocity distribution at the midplane is Maxwellian. This requires that the plasma be deeply (centrifugally) trapped, so that loss regions do not significantly modify the kinetics. It also requires that no other major nonthermal features are present. This assumption would fail, for example, in the presence of a large “sloshing” population of beam ions. The same assumption also takes the plasma to be non-relativistic. At high temperatures (i.e., $T_{e}$ approaching 511 keV), relativistic effects become important [20].
Second, the assumption of quasineutrality is safe only in the core of the plasma. Near the boundaries of the plasma, there are sheath regions in which quasineutrality fails. This is true in both conventional mirrors [21, 22, 23, 24] and centrifugal traps [11, 25, 9]. These sheaths are necessary in order to understand the relationship between the model used in this paper and the treatments of self-consistent electric fields found elsewhere in the literature. It is common to calculate the total voltage drop across the plasma by equating the ion and electron particle loss rates, the idea being that the plasma will charge positive until these rates balance. The sheaths explain how the total voltage drop required by matching loss rates can be consistent with the local voltage profile predicted by quasineutrality. The quasineutrality assumption means that much of the analysis presented here would not apply as written to non-neutral traps like those sometimes used to contain antimatter [26, 27, 28]. However, one can anticipate adapting many of the underlying ideas to those applications – particularly the use of one species to generate fields that trap another species.
@@
The ideas presented here would also apply to forces other than centrifugal forces. Ponderomotive forces have been proposed for various uses in plasma traps [29, 30, 31]; these forces can have various mass and charge dependences, depending on the polarization of the fields. The same analysis would also apply to gravitational forces. Indeed, the mechanisms described here are closely related to effects that have been noted in the astrophysics literature in which protons can be expelled from stars, depending on the stars’ compositions [32, 33]. However, the most immediate application is a new way of end-plugging centrifugal mirror machines. Centrifugal mirrors have advantages over their conventional counterparts, both in terms of confinement and in terms of stability [12, 11, 34, 35, 36, 37]. If the techniques proposed in this paper can further improve the performance of centrifugal mirror traps, the upside potential could be significant.
@@ Appendix A Appendix: More General Distributions
Suppose particles are lost if they get past some maximal energetic barrier height $\varepsilon_{\text{max}}$. Then we can model the distribution as a “truncated Maxwellian,” with

### v2

# Ion Mix Can Invert Centrifugal Confinement

@@ Abstract
Centrifugal plasma traps, in which plasma is confined partly by centrifugal forces, represent a possible path to fusion energy production. In centrifugal plasma traps, electric fields naturally arise in the direction parallel to the magnetic field in order to ensure quasineutrality. These electric fields are sensitive to the particular mix of ion species present. This paper uses analytic and numerical calculations to consider how these mix effects could be used to improve device performance. Changing the composition of the plasma can mix or demix different ion species, improve confinement of certain ions, and can even expel certain species from the trap altogether. This makes it possible to generate an “inverted centrifugal” end-plug with very promising properties.
Introduction. In plasma traps of various kinds, electric fields arise to enforce quasineutrality. These fields are sometimes called ambipolar fields. In magnetized plasmas, ambipolar fields are oriented parallel to the magnetic field lines. They are particularly important in linear devices like magnetic mirrors and centrifugal traps [1, 2]. Linear traps have been the subject of increasing interest in recent years, both for nuclear fusion and for other applications (such as nuclear waste processing and mass separation) [3, 4, 5, 6, 7, 8, 9, 10].
@@
These ideas were developed for conventional (that is, non-centrifugal) mirrors. For centrifugal mirrors, the possibilities associated with these ambipolar fields have been relatively understudied. Centrifugal traps – sometimes also called centrifugal mirrors or rotating mirrors – achieve parallel confinement using centrifugal forces. If the magnetic field lines are arranged so that they bend radially inward near the ends of the trap, the centrifugal forces produce a confining effect along the field lines. Rotation can be driven by superimposing approximately-axial magnetic fields with approximately-radial electric fields; the electric fields can be driven by electrodes, neutral beams, or wave-particle interactions [16, 17, 18, 19, 20, 9]. In this paper, we will consider the effects of ion species mix on ambipolar potentials in centrifugal traps. These mix effects lead to surprising outcomes. Adding more of one species can confine or deconfine other populations, or can lead to stratification effects in which different species are localized at different locations along field lines. Even more surprising is that mix effects can also be used to establish a highly advantageous centrifugal end-plug, analogous to the end-plugs used in conventional mirrors but relying on a different set of physical mechanisms. These effects suggest exciting new possibilities for centrifugal mirror fusion, among other applications.
@@
Here $\Delta\phi$ is the electrical potential difference, $\Delta\varphi_{c}$ is the centrifugal potential energy as it would be experienced by a proton, $q_{s}$ is the charge, $\mu_{s}$ is the particle mass (normalized by the proton mass), and we have taken the temperature $T_{s}$ to be constant along the field line. $n_{s0}$ is a normalization constant. $\Delta\phi$ and $\Delta\varphi_{c}$ are both defined relative to some reference point on the field line. In the quasineutral limit,
@@
In this limit, $\Delta\phi$ is set entirely by species $a$. A well in $\Delta\varphi_{c}$ always confines (attracts) species $a$ and a barrier in $\Delta\varphi_{c}$ always repels species $a$; the electrostatic field simply acts to reduce the efficacy of these wells (local minima) or barriers (maxima) by a factor of $1+Z_{a}(T_{e}/T_{a})$.
@@
Figure 1: Different distributions of protons and boron-11 along a field line, with field line position parameterized by centrifugal potential $\Delta\varphi_{c}$ (normalized by the temperature $T$). As the relative density of the protons drops, the self-consistent ambipolar fields push them further outwards (the top panel has the highest boron concentration; the bottom panel has the highest proton concentration). This example takes the ion and electron temperatures to be equal. Note that the overall normalization is arbitrary, so we could imagine the different cases taking place at fixed total ion charge or fixed total ion number. $n_{p0}$ and $n_{B0}$ are defined as the proton and boron densities at $\Delta\varphi_{c}=0$.
@@
Consider, then, another example: a deuterium-tritium plasma that is “doped” with a smaller fraction of protons. Protons, with their high charge-to-mass ratio, tend to reduce the ambipolar fields wherever they are found. This is illustrated in Fig. 2, where the introduction of a small population of protons substantially improves the confinement of deuterium and tritium in a centrifugal well. The example in the figure introduces a relatively small proton fraction; the effect on deuterium and tritium confinement depends on the proton fraction per Eqs. (1) and (2).
So, should deuterium-tritium centrifugal traps operate with a small fraction of protons? The answer depends on which factors limit the performance of the trap. The introduction of protons allows the deuterium and tritium to be confined at lower rotation speeds, which would suggest that a pressure-limited device could operate at higher plasma densities (the limits on the plasma pressure are less stringent if the plasma rotates more slowly, since the magnetic field does not need to counteract such large centrifugal forces [22]). Note that there will be some loss of fuel density due to dilution; in the limiting case where all of the fuel has been replaced by protons, the reactivity vanishes altogether. In favorable scenarios (particularly when the proton fraction is not too large) this can be more than offset by the increase in overall plasma density; see Fig. 2. However, the protons themselves are relatively poorly confined in these scenarios. This downside might be mitigated using the centrifugal end-plugs proposed later in this paper.
@@
The possibility of better parallel confinement addresses one of the major challenges facing centrifugal traps: strong centrifugal requirement requires fast rotation, and fast rotation requires imposing large electric fields across the magnetic flux surfaces. If the same confinement can be achieved with smaller electric fields, then all of the technical challenges associated with these large fields become less severe.
The same physical mechanisms discussed in the two examples above can also be applied to a wide variety of other problems. One of these is the physics of heavy impurities. Because heavy impurities are often not fully ionized, their charge-to-mass ratio can be particularly low. This means that impurity accumulation has the potential to worsen the longitudinal confinement of all other species. Another straightforward application is plasma mass filters, which are designed to separate an input of mixed ions according to species [23, 24, 25, 26, 27, 28, 5, 29]. If the composition of a centrifugal trap can be tuned to confine or deconfine different species according to mass and charge, the configuration could be used for separations. Centrifugal potentials can also be important in toroidal devices like tokamaks. Distributions of different species along field lines in toroidal devices are often studied using the same analytic models employed here [30, 31, 32], and many of the ideas suggested here could readily be applied to those systems.
@@
This follows from Eq. (5). The confinement generally gets worse if other ions are placed in the well alongside the protons. However, suppose instead that the protons are placed in a region between two centrifugal wells, each populated with some other, lower charge-to-mass ratio ion – say, 7Li. In the limit where the centrifugal wells are populated entirely with lithium, the protons are repelled from the wells, seeing a total potential
@@
In other words, counter-intuitively, for a given well depth $\Delta\varphi_{c}$, it is more efficient for the purposes of proton confinement to set up a barrier full of some other, lower-$Z/m$ species than it is to use that same $\Delta\varphi_{c}$ as a well to directly trap the protons. This is illustrated in Figure 3; the centrifugal wells in the figure correspond to the end-plugs, with the region between the wells corresponding to the central cell  [^1] See Supplemental Material [URL] for a more quantitative scenario.. In order to easily maintain the different compositions in the different segments of the trap, it would be important to make sure that each species is sufficiently well-trapped within its intended region (as in the case in the figure). In some cases, it might be helpful to ensure that the potential barrier between end-cell and the exterior of the trap is weaker than that between end-cell and central cell (for the end-cell species), so that end-cell ions diffuse out of the trap before entering the central cell. One could make an even more effective trap by placing the protons in a centrifugal well of their own, then flanking that well with end-cells full of another species.
@@
Another possibility is to use partially ionized species in the end-cell. For partially ionized species, the charge-to-mass ratio can be very small, so an end-cell populated by a partially ionized species can in principle produce a very strong barrier. For higher-temperature applications, there is the challenge of how to maintain a lower ionization state (and the radiative losses normally associated with partially ionized species like tungsten in high-temperature plasmas). A degree of thermal isolation between the main cell and the end-cells naturally arises if the end-cells produce potential barriers to repel main-cell ions, but the limits on how easily a temperature difference might be maintained are left as an open question. The “thermal barrier” concept for the end-cells of conventional (non-centrifugal) mirrors does rely on maintaining a temperature difference between end-cell and main-cell, and least for the electron population [11, 2]. However, it is substantially easier to maintain a difference in electron temperatures than in ion temperatures. Of course, in the limit of very low barrier species charge-to-mass ratio, it would be important to maintain a large enough magnetic field in the end-cells to maintain perpendicular confinement.
@@
The model used here comes with a number of assumptions. First, and perhaps most importantly: we take each species to be Gibbs-distributed along field lines. This assumption is examined in further detail in the Appendix. Kinetically, it is equivalent to the assumption that the velocity distribution at the midplane is Maxwellian. This requires that the plasma be deeply (centrifugally) trapped, so that loss regions do not significantly modify the kinetics. It also requires that no other major nonthermal features are present. This assumption would fail, for example, in the presence of a large “sloshing” population of beam ions. However, it is consistent with experimental observations in the literature [34, 32]. The same assumption also takes the plasma to be non-relativistic. At high temperatures (i.e., $T_{e}$ approaching 511 keV), relativistic effects become important [35].
Second, the assumption of quasineutrality is safe only in the core of the plasma. Near the boundaries of the plasma, there are sheath regions in which quasineutrality fails. This is true in both conventional mirrors [1, 36, 37, 38] and centrifugal traps [22, 39, 9]. These sheaths are necessary in order to understand the relationship between the model used in this paper and the treatments of self-consistent electric fields found elsewhere in the literature. It is common to calculate the total voltage drop across the plasma by equating the ion and electron particle loss rates, the idea being that the plasma will charge positive until these rates balance. The sheaths explain how the total voltage drop required by matching loss rates can be consistent with the local voltage profile predicted by quasineutrality. A more detailed treatment of the sheaths could be desirable in order to apply these ideas to devices like fusion thrusters. For example, in a deuterium-tritium rocket, tritium may be much more valuable to retain than deuterium. The quasineutrality assumption means that much of the analysis presented here would not apply as written to non-neutral traps like those sometimes used to contain antimatter [40, 41, 42]. However, one can anticipate adapting many of the underlying ideas to those applications – particularly the use of one species to generate fields that trap another species.
@@
The ideas presented here would also apply to forces other than centrifugal forces. Ponderomotive forces have been proposed for various uses in plasma traps [43, 44, 45]; these forces can have various mass and charge dependences, depending on the polarization of the fields. The same analysis would also apply to gravitational forces. Indeed, the mechanisms described here are closely related to effects that have been noted in the astrophysics literature in which protons can be expelled from stars, depending on the stars’ compositions [46, 47]. Other forces would appear in the Gibbs factor in Eq. (1), consistent with a zero-current force balance. However, the most immediate application is a new way of end-plugging centrifugal mirror machines. Centrifugal mirrors have advantages over their conventional counterparts, both in terms of confinement and in terms of stability [23, 22, 48, 49, 34, 50]. If the techniques proposed in this paper can further improve the performance of centrifugal mirror traps, the upside potential could be significant.
@@
Data Availability Statement
The data that support the findings of this article are openly available at [51].
@@ Appendix A Appendix: More General Distributions
Suppose particles are lost if they get past some maximal energetic barrier height $\varepsilon_{\text{max}}$. Different analytic models for loss-cone distributions in low-collisionality plasmas exist [53], but one straightforward choices is to model the distribution as a “truncated Maxwellian,” with
@@
Note that distributions of the form $f(\varepsilon,\mu)$ implicitly assume a low-collisionality regime where $\mu$-invariance is respected (e.g., over the course of many bounces). The analysis in the main body of this paper would apply whether the plasma is collisional or not. When we discuss non-constant temperatures, in the low-collisionality limit we typically have in mind multiple potential wells where the trapped populations in each well may have different characteristic energies. It is also worth pointing out that even in the “collisionless” limit (where the collision time is long compared with the bounce time), collisions may still encourage $f$ to be more nearly Maxwellian, if the collision time is short compared with the confinement time. This is part of the reason why it makes sense to focus on near-Maxwellian models.